# A Framework for Evaluating Emerging Cyberattack Capabilities of AI

### v2

@@ Abstract
As frontier AI models become more capable, evaluating their potential to enable cyberattacks is crucial for ensuring the safe development of Artificial General Intelligence (AGI). Current cyber evaluation efforts are often ad-hoc, lacking systematic analysis of attack phases and guidance on targeted defenses. This work introduces a novel evaluation framework that addresses these limitations by: (1) examining the end-to-end attack chain, (2) identifying gaps in AI threat evaluation, and (3) helping defenders prioritize targeted mitigations and conduct AI-enabled adversary emulation for red teaming. Our approach adapts existing cyberattack chain frameworks for AI systems. We analyzed over 12,000 real-world instances of AI use in cyberattacks catalogued by Google’s Threat Intelligence Group. Based on this analysis, we curated seven representative cyberattack chain archetypes and conducted a bottleneck analysis to pinpoint potential AI-driven cost disruptions. Our benchmark comprises 50 new challenges spanning various cyberattack phases. Using this benchmark, we devised targeted cybersecurity model evaluations, report on AI’s potential to amplify offensive capabilities across specific attack phases, and offer recommendations for prioritizing defenses. We believe this represents the most comprehensive AI cyber risk evaluation framework published to date.
@@
We analyzed over 12,000 instances of real-world AI use attempts in cyberattacks across 20+ countries, catalogued by Google’s Threat Intelligence Group. Based on this, we curated 7 representative cyberattack chain archetypes and conducted a bottleneck analysis to identify phases where AI could most likely disrupt costs. Collaborating with external stakeholders, we developed a new AI cyber capability benchmark with 50 challenges across the attack chain, covering intelligence gathering, operational security, vulnerability exploitation, and malware development. These challenges are novel and not based on public ones, mitigating benchmark contamination risks. To our knowledge, this is the most comprehensive benchmark covering offensive cyber capabilities across the entire attack chain.
@@ 5.3 Evaluation Benchmark Details
Following bottleneck identification (Appendix A), we designed evaluations measuring AI’s capacity to overcome these attack impediments. To avoid overfitting to public data, we partnered with Pattern Labs to develop 50 bespoke challenges based on cybersecurity "first principles" rather than existing competitions. These evaluations test diverse capabilities relevant to bottlenecks, covering a spectrum of difficulty.
@@ 6 Evaluation
Model Configuration. We evaluated Gemini 2.0 Flash experimental. The model was provided with scaffolding (tools, prompting procedure, error handling) to form an agent (Figure 9). The default hyperparameters were used. The workflow involved a controller feeding challenge descriptions to the model, handling tool calls, forwarding commands to the environment, receiving results, and iterating until solved or interaction limit reached.
@@ Acknowledgement
We thank Lewis Ho for insightful reviews; Matthew Rahtz and Alex Kaskasoli for Frontier Safety team support; Ivan Petrov for discussions on emerging capabilities; Jennifer Beroshi, Elie Bursztein, Xerxes Dotiwalla, Gena Gibson, Myriam Khan, Armin Senoner, Rohin Shah, Andy Song, and Andreas Terzis for guidance; Pattern Labs and other external partners; Google’s Threat Intelligence Group, and Google for a supportive research environment.
Finally, we thank Alexandru Totolici, Ansh Chandnani, Ash Fox, Daniel Fabian, Fu Chai, Maksim Shudrak, Mónica Carranza, Niru Ragupathy, Ryan Goosen, and Stefan Friedli, for their valuable time and insight into resource bottlenecks for historical cyberattack case studies.
@@ A.2 Validating Attack Bottlenecks via Expert Survey
We define a ’bottleneck’ as any phase requiring at least 10% of the total estimated resources for the attack. Our key findings were that:
@@
Averages across case studies where a given phase was a bottleneck:
Weaponization (n=9, 35.2%)
Reconnaissance (n=7, 19.3%)
Actions on Objectives (n=5, 17.6%)
Exploitation (n=4, 8.3%)
Delivery (n=4, 8.3%)
C2 (n=3, 6.4%)
Installation (n=2, 3.2%)

### v3

@@ Abstract
As frontier AI models become more capable, evaluating their potential to enable cyberattacks is crucial for ensuring the safe development of Artificial General Intelligence (AGI). Current cyber evaluation efforts are often ad-hoc, lacking systematic analysis of attack phases and guidance on targeted defenses. This work introduces a novel evaluation framework that addresses these limitations by: (1) examining the end-to-end attack chain, (2) identifying gaps in AI threat evaluation, and (3) helping defenders prioritize targeted mitigations and conduct AI-enabled adversary emulation for red teaming. Our approach adapts existing cyberattack chain frameworks for AI systems. We analyzed over 12,000 real-world instances of AI involvement in cyber incidents, catalogued by Google’s Threat Intelligence Group, to curate seven representative attack chain archetypes. Through a bottleneck analysis on these archetypes, we pinpointed phases most susceptible to AI-driven disruption. We then identified and utilized externally developed cybersecurity model evaluations focused on these critical phases. We report on AI’s potential to amplify offensive capabilities across specific attack stages, and offer recommendations for prioritizing defenses. We believe this represents the most comprehensive AI cyber risk evaluation framework published to date.
@@
The benchmark is comprised of the following elements:
A curated set of representative cyberattack chain archetypes derived from analyzing over 12,000 instances of real-world AI use attempts in cyberattacks across 20+ countries.
A set of of bottlenecks identified across each of the attack chain archetypes
A set of human expert baselines that can be used to calibrate results and refine the selection of bottlenecks A set of relative weights associated with specific TTPs across each stage of the attack chain based on prevalence in the wild
A set of 50 externally developed CTFs and representative environments sourced from Pattern Labs that are not public and therefore mitigate the risk of training data contamination
@@ 5.3 Evaluation Benchmark Details
Following bottleneck identification (Appendix A), we sourced 50 evaluations from Pattern Labs’ library of withheld CTFs that test diverse capabilities relevant to the bottlenecks we identified, covering a spectrum of difficulty.
@@ 6 Evaluation
Model Configuration. We evaluated Gemini 2.0 Flash experimental. To ensure consistency of the results and findings we conducted our evaluation using both an internal scaffolding (tools, prompting procedure, error handling) to form an agent (Figure 9) as well as using the scaffolding provided by the vendor that developed the CTF challenges.The default hyperparameters were used. The workflow involved a controller feeding challenge descriptions to the model, handling tool calls, forwarding commands to the environment, receiving results, and iterating until solved or interaction limit reached.
@@ Acknowledgement
We thank Lewis Ho for insightful reviews; the larger Frontier Safety team for collaboration on frontier evaluations; Ivan Petrov for discussions on emerging capabilities; Jennifer Beroshi, Elie Bursztein, Xerxes Dotiwalla, Gena Gibson, Myriam Khan, Armin Senoner, Rohin Shah, Andy Song, and Andreas Terzis for guidance; Alexandru Totolici, Ansh Chandnani, Ash Fox, Daniel Fabian, Fu Chai, Maksim Shudrak, Mónica Carranza, Niru Ragupathy, Ryan Goosen, and Stefan Friedli, for their valuable time and insight into resource bottlenecks for historical cyberattack case studies; Pattern Labs specifically to Dan Lahav, Omer Nevo, Ofir Ohad, and Saar Tochner for access to their evaluation platform and library; Google’s Threat Intelligence Group for relevant data and insights; and broadly Google for a supportive research environment.
@@ A.2 Validating Attack Bottlenecks via Expert Survey
We took the initial human-day estimates and converted them to dollar costs by multiplying by a salary variable ($500k); for ease of cost comparison between phases, we then calculated the percent of the overall resourcing estimate each phase accounted for. We define a ’bottleneck’ as any phase requiring at least 10% of the total estimated resources for the attack.
Our key findings were that:
@@
Averages across case studies:
Weaponization (n=9, 35.94%, $206.7k)
Reconnaissance (n=7, 19.20%, $44.8k)
Actions on Objectives (n=5, 17.60%, $525k; 11.00%, $16.5k when excluding SolarWinds outlier)
Exploitation (n=4, 9.82%, $19.3k)
Delivery (n=4, 8.30%, $15.8k)
C2 (n=3, 6.40%, $17.3k)
Installation (n=2, 3.20%, $8.3k)